A number is 27 more than the number obtained by reversing its digits. If its unit's and ten's digit are x and y respectively, write the linear equation representing the above statement.
Answer by ankor@dixie-net.com(22740)

let x = the 10's digit
let y the units
then
10x + y = "A number"
:
A number is 27 more than the number obtained by reversing its digits.
10x + y = 10y + x + 27
10x - x = 10y - y + 27
9x = 9y + 27
simplify, divide thru by 9
x = y + 3